# Answers to frequently asked questions about the pulsar timing array Hellings and Downs curve

### v2

@@ 4 Answers to frequently asked questions
The reason why $1/3$ and $1/2$ have appeared in the literature is historical. The Hellings and Downs curve was first computed as the average product of the redshift responses of two pulsars to a unit-amplitude GW source, averaged over the sphere of GW source directions as described in Section 3.1. This average introduces a factor of $1/4\pi$ in (1cdmatb). Later, the Hellings and Downs curve was reinterpreted as a correlation coefficient for the redshift response of two pulsars, responding to the same GW source. For this case, one wants the correlation coefficient to equal unity if the two pulsars are identical (i.e., $a=b$). Since the expected correlation for two identical pulsars is twice that for two distinct pulsars separated by $0^{\circ}$ [due to the $\delta_{ab}$ term in (1cdmas)], one needs $\mu_{\rm u}(0)=1/2$ for the correlation-coefficient interpretation. This means multiplying the right-hand sides of (1cdmatb) and (1cdmatau) by 3/2.
@@
Answer: To prove this, start with (1cdmatb) for the expected Hellings and Downs correlation and (1cdmu) for the Earth-term-only response functions $F^{A}(\hat{\Omega})$. Taking the two pulsars to point in the same direction ($\hat{p}_{1}=\hat{p}_{2}=\hat{z}$), we have
$\sum_{A}F_{1}^{A}(\hat{\Omega})F_{2}^{A}(\hat{\Omega})=\frac{1}{4}(1+\cos\theta)^{2}=\frac{1}{4}(1+u)^{2}\,,$ (1cdmatay)
@@
$\sum_{A}F_{1}^{A}(\hat{\Omega})F_{2}^{A}(\hat{\Omega})=\frac{1}{4}(1-\cos^{2}\theta)=\frac{1}{4}(1-u^{2})\,.$ (1cdmataz)
@@
$\mu_{\rm u}(0^{\circ})=\frac{1}{3}\quad(\hat{p}_{1}=\hat{p}_{2}=\hat{z})\,,\quad\mu_{\rm u}(180^{\circ})=\frac{1}{6}\quad(\hat{p}_{1}=-\hat{p}_{2}=\hat{z})\,.$ (1cdmatba)
@@
    Why isn’t the most negative value of the Hellings and Downs curve at $90^{\circ}$?
Answer: Again, this is because the Earth-term-only response functions $F^{A}(\hat{\Omega})$ given in (1cdmu) are not symmetric under $\hat{p}\rightarrow-\hat{p}$. If the Hellings and Downs correlation were purely quadrupolar, then its most negative value would be exactly at $90^{\circ}$, as it is for the case of an array of short-arm LIGO-like detectors (see the right panel of Figure 9). If it had only even multipoles (no odd-$l$ terms) then it would be reflection symmetric about $90^{\circ}$ but might not have its minimum there.
@@
$\mu_{\rm u}(\gamma)=\sum_{l=2}^{\infty}(2\ell+1)c_{l}P_{l}(\cos\gamma)\,,\quad{\rm where}$ (1cdmatbb)
$c_{l}=[(l+2)(l+1)l(l-1)]^{-1}\,.$ (1cdmatbc)
@@
$\int_{-1}^{1}{\rm d}x\>P_{l}(x)P_{l^{\prime}}(x)=\frac{2}{2l+1}\delta_{ll^{\prime}}$ (1cdmatbd)
of the Legendre polynomials, where $x\equiv\cos\gamma$, (b) Rodrigues’ Formula for Legendre polynomials, and (c) integration by parts. See equations (97)–(105) in [12] for a step-by-step derivation of (1cdmatbb) and (1cdmatbc).
Figure 11 compares approximate versions of the Hellings and Downs curve, terminating the sum at different maximum values $l_{\rm max}=2,3,4,5$. Since the $c_{l}$’s in (1cdmatbc) fall off like $1/l^{4}$, only a few values of $l$ are needed to get a very good approximation.
Figure 11: Approximations to the Hellings and Downs curve (solid black) obtained by terminating the sum in (1cdmatbc) with two to five Legendre polynomials (grey curves).
@@
$R_{a}^{A}(f,\hat{\Omega})=F^{A}(\hat{\Omega})e^{-i2\pi f\hat{\Omega}\cdot\vec{r}_{a}}\,,$ (1cdmatbea)
$F_{a}^{A}(\hat{\Omega})\equiv\frac{1}{2}(\hat{X}_{a}^{i}\hat{X}_{a}^{j}-\hat{Y}_{a}^{i}\hat{Y}_{a}^{j})e^{A}_{ij}(\hat{\Omega})\,,\quad A=+,\times\,,$ (1cdmatbeb)
@@
$\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>\sum_{A}F^{A}_{a}(\hat{\Omega})F^{A}_{b}(\hat{\Omega})\,e^{-i2\pi f\hat{\Omega}\cdot(\vec{r}_{a}-\vec{r}_{b})}\,.$ (1cdmatbebf)
@@
As discussed in Section 3.1, the original/traditional approach to compute the expected Hellings and Downs correlation is to take a single fixed pair of pulsars, and then to average their correlation over GW propagation directions $\hat{\Omega}$, assuming that the sources are unpolarized and isotropic on the sky [5]. In references [18] and [9], somewhat involved calculations were done to show the equivalence of pulsar averaging and source averaging for the case of a single GW source, see (1cdmatax).
Happily, one can show the equivalence of these two approaches without any calculation. Here, we show that pulsar averaging is equivalent to GW-source averaging, with a purely geometric argument, using a series of three Euler-angle rotations. At first, this appears to require adding something to the traditional source-averaging calculation: in addition to averaging over the GW propagation directions $\hat{\Omega}=\hat{\Omega}(\theta,\phi)$ one must also average over the polarization angle $\psi$. In fact, while often overlooked, this is also required by the traditional calculation: it gives rise to the quadrature sum. It is overlooked because the original [5] calculation obscures this point by not giving enough detail. To obtain that result, one must assign a polarization direction or angle to each source, and also average over that, for example, as in the calculation starting from (1cdmac). Hence, even in the traditional GW-source averaging approach, the averages must be calculated with respect to three angles, not just two.
To see the equivalence purely from geometry, start with the averaging over GW source directions. As is often done, without loss of generality, put $\hat{p}_{1}$ along the $z$-axis and $\hat{p}_{2}$ in the $xz$-plane, so that $\hat{p}_{2}=\sin\gamma\,\hat{x}+\cos\gamma\,\hat{z}$. In contrast, pulsar averaging is carried out with respect to the angular coordinates of $\hat{p}_{1}$ (polar coordinate $\theta$, azimuthal coordinate $\psi$) and those of $\hat{p}_{2}$ (which lies on a cone about $\hat{p}_{1}$ with fixed half-opening angle $\gamma$, with its location on the cone parameterized by $\phi\in[0,2\pi)$). See Figure 12 and compare with Figure 5 of [9].
@@
If we have an infinite collection of universes, each of which has GW sources having independent random phases, and we were also able to average the correlation over all of those universes (called an “ensemble average”), then we would get exactly the Hellings and Downs curve. But in any given universe, we would find something different from Hellings and Downs. The (mean squared) difference between what would be observed in any given universe, and the (unobservable) average over many universes, is called the cosmic variance.
It is easy to see how the cosmic variance arises from interference between different GW sources. Consider a universe which contains $N$ point sources of GWs, which we label $n=1,\dots,N$. These GW sources are located far from the neighborhood containing the Earth and the PTA pulsars. The GW sources have sky locations $-\hat{\Omega}_{n}$, and the $n$’th source produces a complex strain at Earth $h_{ij}(t)=\Re[h_{n}(t)e^{*}_{ij}(\hat{\Omega}_{n})]$, where the complex waveform is $h_{n}(t)=h_{n}^{+}(t)+ih_{n}^{+}(t)$. Since the PTA is far from the sources, these GW strains are small, so the total strain at Earth can be obtained by summing them. The redshift of pulsar $a$ is then obtained by summing (1cdms) to obtain
$Z_{a}(t)$ $=\Re\left[\sum_{n=1}^{N}h_{n}(t)F^{*}_{a}(\hat{\Omega}_{n})\right]$ (1cdmatbebga)
$=\frac{1}{2}\sum_{n}\Bigl{[}h_{n}(t)F^{*}_{a}(\hat{\Omega}_{n})+h^{*}_{n}(t)F_{a}(\hat{\Omega}_{n})\Bigr{]}\,,$ (1cdmatbebgb)
where $F_{a}(\hat{\Omega})\equiv F_{a}^{+}(\hat{\Omega})+iF_{a}^{\times}(\hat{\Omega})$ are the complex antenna patterns, and we adopt the convention that sums over $n$ and $m$ are from $1,\cdots,N$ unless otherwise indicated. A similar expression holds for pulsar $b$. We assume that the pulsars $a$ and $b$ are distinct and are separated from one another by many GW wavelengths, and therefore have neglected the pulsar terms for the reasons discussed earlier. (We work with complex quantities because it simplifies some of the following.)
@@
$\hskip 46.97505pt+\overline{h_{n}h_{m}}F^{*}_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{m})+\overline{h^{*}_{n}h^{*}_{m}}F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{m})\Bigr{]}\,.$ (1cdmatbebgbh)
@@
Next, we consider the pulsar average of (1cdmatbebgbh), which is obtained by averaging the correlation over all pulsar pairs $ab$ separated by angle $\gamma$. As discussed briefly in Sec. 3.2, and in more detail in [19], this corresponds closely to experimental practice: PTA experiments “reconstruct” the angular dependence of the Hellings and Downs correlation via a pulsar averaging procedure. Also, note that we dropped the pulsar terms from (1cdmatbebga). Had we included them, they would have been eliminated by this pulsar averaging step, as shown in Sec. IV of [9].
Consider first the case where there is no interference between GW sources, for example, because the sources are radiating GWs at different frequencies. This means that all the time averages $\overline{h_{n}h_{m}}$ and $\overline{h_{n}h^{*}_{m}}$ vanish if $n$ is different than $m$. Thus, in this no-interference case, the double sum reduces to a single sum:
@@
$\hskip 119.24506pt+2\Re\left(F^{*}_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{n})\,\overline{h^{2}_{n}}\right)\Big{]}.$ (1cdmatbebgbi)
We now pulsar-average the correlation (1cdmatbebgbi). The only dependence on $a$ and $b$ in (1cdmatbebgbi) is via the antenna pattern functions. As explained in Q6 and Figure 12, the pulsar average of $F_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{n})$ is independent of the source direction $\hat{\Omega}_{n}$ and gives exactly the Hellings and Downs curve $\mu_{\rm u}(\gamma)$. The pulsar average of $F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{n})$ vanishes. Hence, the pulsar average of (1cdmatbebgbi) gives
$\langle\rho_{ab}\rangle_{ab\in\gamma}=\frac{1}{2}\Bigl{(}\sum_{n}\overline{|h_{n}|^{2}}\Bigr{)}\mu_{\rm u}(\gamma)\,,$ (1cdmatbebgbj)
where the subscript “$ab\in\gamma$” is a reminder that we have averaged over pulsar pairs separated by angle $\gamma$. Thus, we have proved that if there is no interference between the GW sources, then the pulsar average of the correlation exactly matches the Hellings and Downs curve in any instance of the universe.
@@
If the GW sources interfere, then the pulsar average of (1cdmatbebgbh) contains the “diagonal” $n=m$ terms whose pulsar average we found in (1cdmatbebgbj), and also has additional contributions arising from the off-diagonal $n\neq m$ interference terms. To understand the effects of these interference terms, it is helpful to introduce a real function of two variables $\mu(\gamma,\beta)$. This was first defined and computed by Allen in Appendix G of [9], who named it the “Hellings and Downs two-point function”. It is given explicitly in Eq. (G5) of [9] and plotted in Figure 12 of that reference. We will assume that the reader has looked at these, and used Eqs. (G9) and (G10) of [9] to verify that the pulsar average of $F_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{m})$ is $\mu(\gamma,\beta_{nm})$ and that the pulsar average of $F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{m})$ is $\mu(\pi-\gamma,\pi-\beta_{nm})$. Here, $\gamma=\gamma_{ab}$ is the angle between the directions to the two pulsars ($\cos\gamma=\hat{p}_{a}\cdot\hat{p}_{b}$), and $\beta_{nm}$ is the angle between the directions to the two GW sources ($\cos\beta_{nm}=\hat{\Omega}_{n}\cdot\hat{\Omega}_{m}$). Note that in the case where the two GW sources are the same, $\beta_{nm}=0$ and the two-point function reduces to the normal Hellings and Downs curve: $\mu(\gamma,0)=\mu_{\rm u}(\gamma)$.
@@
Thus, in the case where the GW sources interfere, the pulsar average of (1cdmatbebgbh) gives
$\hskip 26.73953pt\langle\rho_{ab}\rangle_{ab\in\gamma}=\frac{1}{2}\left(\sum_{n}\overline{|h_{n}|^{2}}\right)\mu_{\rm u}(\gamma)$
$+\frac{1}{2}\sum_{n\neq m}\left[\Re\left(\overline{h_{n}h^{*}_{m}}\right)\mu(\gamma,\beta_{nm})+\Re\left(\overline{h_{n}h_{m}}\right)\mu(\pi-\gamma,\pi-\beta_{nm})\right]\,.$ (1cdmatbebgbk)
The key point is this: the additional $N(N-1)$ terms arising from the two-point function, considered as functions of the angle $\gamma$, are not proportional to the Hellings and Downs curve $\mu_{\rm u}(\gamma)$. Moreover, even when the number of sources is large, these terms make contributions to the pulsar average which are comparable to the first (diagonal) term.
A simple explicit example is the “confusion-noise model” of [9]. Each of the $N$ sources of GWs is circularly polarized, and all of them radiate at the same frequency, which is picked to be commensurate with the PTA observation interval. From (1cdi), the waveforms are
$h_{n}(t)=A_{n}\exp[i(2\pi ft+\phi_{n})]\,,$ (1cdmatbebgbl)
where the amplitudes $A_{n}>0$ are real, the GW frequency $f$ is the same fixed multiple of $1/T$ for all $N$ sources, and the phase $\phi_{n}$ of the $n$’th source is a real number in the interval $[0,2\pi)$. For this model, the time averages which appear above are given by $\overline{h_{n}h_{m}}=0$ and $\overline{h_{n}h^{*}_{m}}=A_{n}A_{m}\exp[i(\phi_{n}-\phi_{m})]$. Thus, the pulsar average in (1cdmatbebgbk) becomes
@@
$\hskip 36.135pt\langle\rho_{ab}\rangle_{ab\in\gamma}=\frac{1}{2}\Bigl{(}\sum_{n}A^{2}_{n}\Bigr{)}\mu_{\rm u}(\gamma)+\frac{1}{2}\sum_{n\neq m}A_{n}A_{m}\cos(\phi_{n}-\phi_{m})\,\mu(\gamma,\beta_{nm})\,.$

### v3

@@ 4 Answers to frequently asked questions
The reason why $1/3$ and $1/2$ have appeared in the literature is historical. The Hellings and Downs curve was first computed as the average product of the redshift responses of two pulsars to a unit-amplitude GW source, averaged over the sphere of GW source directions as described in Section 3.1. This average introduces a factor of $1/4\pi$ in (1cdmaub). Later, the Hellings and Downs curve was reinterpreted as a correlation coefficient for the redshift response of two pulsars, responding to the same GW source. For this case, one wants the correlation coefficient to equal unity if the two pulsars are identical (i.e., $a=b$). Since the expected correlation for two identical pulsars is twice that for two distinct pulsars separated by $0^{\circ}$ [due to the $\delta_{ab}$ term in (1cdmat)], one needs $\mu_{\rm u}(0)=1/2$ for the correlation-coefficient interpretation. This means multiplying the right-hand sides of (1cdmaub) and (1cdmauav) by 3/2.
@@
Answer: To prove this, start with (1cdmaub) for the expected Hellings and Downs correlation and (1cdmu) for the Earth-term-only response functions $F^{A}(\hat{\Omega})$. Taking the two pulsars to point in the same direction ($\hat{p}_{1}=\hat{p}_{2}=\hat{z}$), we have
$\sum_{A}F_{1}^{A}(\hat{\Omega})F_{2}^{A}(\hat{\Omega})=\frac{1}{4}(1+\cos\theta)^{2}=\frac{1}{4}(1+u)^{2}\,,$ (1cdmauaz)
@@
$\sum_{A}F_{1}^{A}(\hat{\Omega})F_{2}^{A}(\hat{\Omega})=\frac{1}{4}(1-\cos^{2}\theta)=\frac{1}{4}(1-u^{2})\,.$ (1cdmauba)
@@
$\mu_{\rm u}(0^{\circ})=\frac{1}{3}\quad(\hat{p}_{1}=\hat{p}_{2}=\hat{z})\,,\quad\mu_{\rm u}(180^{\circ})=\frac{1}{6}\quad(\hat{p}_{1}=-\hat{p}_{2}=\hat{z})\,.$ (1cdmaubb)
@@
    Why is the most negative value of the Hellings and Downs curve not at $90^{\circ}$?
Answer: Again, this is because the Earth-term-only response functions $F^{A}(\hat{\Omega})$ given in (1cdmu) are not symmetric under $\hat{p}\rightarrow-\hat{p}$. If the Hellings and Downs correlation were purely quadrupolar, then its most negative value would be exactly at $90^{\circ}$, as it is for the case of an array of short-arm LIGO-like detectors (see the right panel of Figure 9). If it had only even multipoles (no odd-$l$ terms), then it would be reflection symmetric about $90^{\circ}$, but it might not have its minimum there.
@@
$\mu_{\rm u}(\gamma)=\sum_{l=2}^{\infty}(2l+1)c_{l}P_{l}(\cos\gamma)\,,\quad{\rm where}$ (1cdmaubc)
$c_{l}=[(l+2)(l+1)l(l-1)]^{-1}\,.$ (1cdmaubd)
@@
$\int_{-1}^{1}{\rm d}x\>P_{l}(x)P_{l^{\prime}}(x)=\frac{2}{2l+1}\delta_{ll^{\prime}}$ (1cdmaube)
of the Legendre polynomials, where $x\equiv\cos\gamma$, (b) Rodrigues’ Formula for Legendre polynomials, and (c) integration by parts. See equations (97)–(105) in [12] for a step-by-step derivation of (1cdmaubc) and (1cdmaubd), or see Secs. III and IV of [17].
Figure 11 compares approximate versions of the Hellings and Downs curve, terminating the sum at different maximum values $l_{\rm max}=2,3,4,5$. Since the $c_{l}$’s in (1cdmaubd) fall off like $1/l^{4}$, only a few values of $l$ are needed to get a very good approximation.
Figure 11: Approximations to the Hellings and Downs curve (solid black) obtained by terminating the sum in (1cdmaubd) with two to five Legendre polynomials (grey curves).
@@
$R_{a}^{A}(f,\hat{\Omega})=F_{a}^{A}(\hat{\Omega})e^{-i2\pi f\hat{\Omega}\cdot\vec{r}_{a}}\,,$ (1cdmaubfa)
$F_{a}^{A}(\hat{\Omega})\equiv\frac{1}{2}(\hat{X}_{a}^{i}\hat{X}_{a}^{j}-\hat{Y}_{a}^{i}\hat{Y}_{a}^{j})e^{A}_{ij}(\hat{\Omega})\,,\quad A=+,\times\,,$ (1cdmaubfb)
@@
$\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>\sum_{A}F^{A}_{a}(\hat{\Omega})F^{A}_{b}(\hat{\Omega})\,e^{-i2\pi f\hat{\Omega}\cdot(\vec{r}_{a}-\vec{r}_{b})}\,.$ (1cdmaubfbg)
@@
As discussed in Section 3.1, the original/traditional approach to compute the expected Hellings and Downs correlation is to take a single fixed pair of pulsars, and then to average their correlation over GW propagation directions $\hat{\Omega}$, assuming that the sources are unpolarized and isotropic on the sky [6], and that they do not interfere with each other. In references [22] and [5], somewhat involved calculations were done to show the equivalence of pulsar averaging and source averaging for the case of a single GW source, see (1cdmauay).
Happily, one can show the equivalence of these two approaches without any calculation. Here, we show that pulsar averaging is equivalent to GW-source averaging with a purely geometric argument, using a series of three Euler-angle rotations. At first, this appears to require adding something to the traditional source-averaging calculation: in addition to averaging over the GW propagation direction $\hat{\Omega}=\hat{\Omega}(\theta,\phi)$, one must also average over the polarization angle $\psi$. In fact, while often overlooked, this is also required by the traditional calculation: it gives rise to the quadrature sum. It is overlooked because the original [6] calculation obscures this point by not giving enough detail. To obtain that result, one must assign a polarization direction or angle to each source, and also average over that, for example, as in the calculation starting from (1cdmac). Hence, even in the traditional GW-source averaging approach, the averages must be calculated with respect to three angles, not just two.
To see the equivalence purely from geometry, start with the averaging over GW source directions. As previously done, without loss of generality, put $\hat{p}_{1}$ along the $z$-axis and $\hat{p}_{2}$ in the $xz$-plane, so that $\hat{p}_{2}=\sin\gamma\,\hat{x}+\cos\gamma\,\hat{z}$. In contrast, pulsar averaging is carried out with respect to the angular coordinates of $\hat{p}_{1}$ (polar coordinate $\theta$, azimuthal coordinate $\psi$) and those of $\hat{p}_{2}$ (which lies on a cone about $\hat{p}_{1}$ with fixed half-opening angle $\gamma$, with its location on the cone parameterized by $\phi\in[0,2\pi)$). See Figure 12 and compare with Figure 5 of [5].
@@
If we have an infinite collection of universes, each of which has GW sources having independent random phases, and we were also able to average the correlation over all of those universes (called an “ensemble average”), then we would get exactly the Hellings and Downs curve. But in any given universe, we would find something different from Hellings and Downs. The (mean squared) difference between what would be observed in any given universe, and the (unobservable) average over many universes, is called the cosmic variance. In what follows, we show how the cosmic variance arises from interference between different GW sources.
Consider a universe which contains $N$ point sources of GWs, which we label $n=1,\dots,N$. These GW sources are located far from the neighborhood containing the Earth and the PTA pulsars. The GW sources have sky locations $-\hat{\Omega}_{n}$, and the $n$’th source produces a complex strain at Earth $h_{ij}(t)=\Re[h_{n}(t)e^{*}_{ij}(\hat{\Omega}_{n})]$, where the complex waveform is $h_{n}(t)=h_{n}^{+}(t)+ih_{n}^{+}(t)$. Since the PTA is far from the sources, these GW strains are small, so the total strain at Earth can be obtained by summing them. The redshift of pulsar $a$ is then obtained by summing (1cdms) to obtain
$Z_{a}(t)$ $=\Re\left[\sum_{n=1}^{N}h_{n}(t)F^{*}_{a}(\hat{\Omega}_{n})\right]$ (1cdmaubfbha)
$=\frac{1}{2}\sum_{n}\Bigl{[}h_{n}(t)F^{*}_{a}(\hat{\Omega}_{n})+h^{*}_{n}(t)F_{a}(\hat{\Omega}_{n})\Bigr{]}\,,$ (1cdmaubfbhb)
where $F_{a}(\hat{\Omega})\equiv F_{a}^{+}(\hat{\Omega})+iF_{a}^{\times}(\hat{\Omega})$ are the complex antenna patterns, and we adopt the convention that sums over $n$ and $m$ are from $1,\cdots,N$ unless otherwise indicated. A similar expression holds for pulsar $b$. We assume that the pulsars $a$ and $b$ are distinct and are separated from one another by many GW wavelengths, and therefore have neglected the pulsar terms for the reasons discussed earlier. (We work with complex quantities because it simplifies some of the following expressions.)
@@
$\hskip 46.97505pt+\overline{h_{n}h_{m}}F^{*}_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{m})+\overline{h^{*}_{n}h^{*}_{m}}F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{m})\Bigr{]}\,.$ (1cdmaubfbhbi)
@@
Consider the pulsar average of (1cdmaubfbhbi), which is obtained by averaging the correlation over all pulsar pairs $ab$ separated by angle $\gamma$. As discussed briefly in Sec. 3.2, and in more detail in [23], this corresponds closely to experimental practice: as illustrated in figure 1, PTA experiments “reconstruct” the angular dependence of the Hellings and Downs correlation via a pulsar averaging procedure. Also, note that we dropped the pulsar terms from (1cdmaubfbha). Had we included them, they would have been eliminated by this pulsar averaging step, as shown in Sec. IV of [5].
Consider first the case where there is no interference between GW sources, for example, because the sources are radiating GWs at different frequencies. This means that the time averages $\overline{h_{n}h_{m}}$ and $\overline{h_{n}h^{*}_{m}}$ vanish if $n$ is different than $m$. Thus, in this no-interference case, the double sum reduces to a single sum:
@@
$\hskip 119.24506pt+2\Re\left(F^{*}_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{n})\,\overline{h^{2}_{n}}\right)\Big{]}.$ (1cdmaubfbhbj)
We now pulsar-average the correlation (1cdmaubfbhbj). The only dependence on $a$ and $b$ in (1cdmaubfbhbj) is via the antenna pattern functions. As explained in Q6 and Figure 12, the pulsar average of $F_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{n})$ is independent of the source direction $\hat{\Omega}_{n}$ and gives exactly the Hellings and Downs curve $\mu_{\rm u}(\gamma)$. The pulsar average of $F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{n})$ vanishes. Hence, the pulsar average of (1cdmaubfbhbj) gives
$\langle\rho_{ab}\rangle_{ab\in\gamma}=\frac{1}{2}\Bigl{(}\sum_{n}\overline{|h_{n}|^{2}}\Bigr{)}\mu_{\rm u}(\gamma)\,,$ (1cdmaubfbhbk)
where the subscript “$ab\in\gamma$” indicates that we have averaged over pulsar pairs separated by angle $\gamma$. Thus, we have proved that if there is no interference between the GW sources, then the pulsar average of the correlation exactly matches the Hellings and Downs curve in any instance of the universe.
@@
If the GW sources interfere, then the pulsar average of (1cdmaubfbhbi) contains the “diagonal” $n=m$ terms whose pulsar average we found in (1cdmaubfbhbk), and also has additional contributions arising from the off-diagonal $n\neq m$ interference terms. To understand the effects of these interference terms, it is helpful to introduce the “Hellings and Downs two-point function”. This was first defined and computed for a particular pair of GW source directions in Appendix G of [5]. It was then generalized to arbitrary source directions in [17], giving rise to an additional unit-modulus complex phase, whose argument $2\chi$ is defined by [17, Eq. (6.5)].
The Hellings and Downs two-point function is the pulsar average of $F_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{m})$ over all pulsar pairs $a$ and $b$, uniformly distributed over the sphere, and separated by angle $\gamma$, as discussed in Sec. 3.2. It may be written in the form
$\langle F_{a}(\hat{\Omega}_{n})F^{*}_{b}(\hat{\Omega}_{m})\rangle_{ab\in\gamma}=\mu(\gamma,\beta_{nm}){\rm e}^{2i\chi^{\phantom{{}^{\prime}}}_{nm}}\,,$ (1cdmaubfbhbl)
where $\mu(\gamma,\beta)$ is a real function of two variables given explicitly in Eq. (G5) of [5] and plotted in Figure 12 of that reference. We will assume that the reader is familiar with this plot; $\gamma$ is the angle between the directions to the two pulsars in the averaging ($\cos\gamma=\hat{p}_{a}\cdot\hat{p}_{b}$), and $\beta_{nm}=\beta_{mn}$ is the angle between the directions to the two GW sources ($\cos\beta_{nm}=\hat{\Omega}_{n}\cdot\hat{\Omega}_{m}$). The real angle $\chi^{\phantom{{}^{\prime}}}_{nm}\equiv\chi(\hat{\Omega}_{n},\hat{\Omega}_{m})$, where $\chi(\hat{\Omega}_{n},\hat{\Omega}_{m})$ is defined by [17, Eq. (6.5)]. Note that $\chi$ is an antisymmetric function of the positions $\hat{\Omega}_{n}$ and $\hat{\Omega}_{m}$ of the sources [^9] While this analysis only needs the symmetry properties of $\chi_{nm}$ and $\chi^{\prime}_{nm}$, for completeness we note that $\chi^{\phantom{{}^{\prime}}}_{nm}\equiv\chi(\hat{\Omega}_{n},\hat{\Omega}_{m})$ and $\chi^{\prime}_{nm}\equiv\chi(\hat{\Omega}_{n},-\hat{\Omega}_{m})$, where $\chi(\hat{\Omega},\hat{\Omega}^{\prime})$ is defined by [17, Eq. (6.5)]. In this definition of $\chi^{\prime}$, the second argument has been shifted to the antipodal point. Thus, if $\hat{\Omega}_{m}$ has spherical polar coordinates $(\theta_{m},\phi_{m})$ then its antipodal point $-\hat{\Omega}_{m}$ has coordinates $(\pi-\theta_{m},\phi_{m}+\pi)$. $\chi^{\phantom{{}^{\prime}}}_{nm}=-\chi^{\phantom{{}^{\prime}}}_{mn}$; it vanishes for the particular GW source locations $\hat{\Omega}$ and $\hat{\Omega}^{\prime}$ considered in [5]. Note that if the two GW sources are coincident, then $\beta_{nm}=\chi^{\phantom{{}^{\prime}}}_{nm}=0$, and the two-point function reduces to $\mu(\gamma,0)=\mu_{\rm u}(\gamma)$, which is the Hellings and Downs curve.
The pulsar average of $F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{m})$ can be obtained from (1cdmaubfbhbl) by sending both $b$ and $\hat{\Omega}_{m}$ to their antipodal points on the sphere, as detailed in [5, 17]. This sends $\gamma\to\pi-\gamma$ and $\beta\to\pi-\beta$ and also modifies the phase; one obtainsfootnotemark:
$\langle F_{a}(\hat{\Omega}_{n})F_{b}(\hat{\Omega}_{m})\rangle_{ab\in\gamma}=\mu(\pi-\gamma,\pi-\beta_{nm}){\rm e}^{2i\chi^{\prime}_{nm}}\,,$ (1cdmaubfbhbm)
where $\chi^{\prime}_{nm}\equiv\chi(\hat{\Omega}_{n},-\hat{\Omega}_{m})$ is real and symmetric, so $\chi^{\prime}_{nm}=\chi^{\prime}_{mn}$; see [17, Eq. (C4)]. If the two GW sources are antipodal, then the two-point function vanishes: $\mu(\gamma,\pi)=0$, as can be seen in Figure 12 of [5].
@@
Thus, in the case where the GW sources interfere, the pulsar average of (1cdmaubfbhbi) gives
$\langle\rho_{ab}\rangle_{ab\in\gamma}$ $=$ $\frac{1}{2}\Bigr{(}\sum_{n}\overline{|h_{n}|^{2}}\Bigr{)}\mu_{\rm u}(\gamma)+$
@@
$\frac{1}{2}\Re\sum_{n\neq m}\Bigl{[}\bigl{(}\overline{h^{*}_{n}h_{m}}{\rm e}^{2i\chi^{\phantom{{}^{\prime}}}_{nm}}\!\bigr{)}\mu(\gamma,\beta_{nm})+\bigl{(}\overline{h^{*}_{n}h^{*}_{m}}{\rm e}^{2i\chi^{\prime}_{nm}}\!\bigr{)}\mu(\pi-\gamma,\pi-\beta_{nm})\Bigr{]}\,.$
The key point is this: the additional $N(N-1)$ terms arising from the two-point function, considered as functions of the angle $\gamma$, are not proportional to the Hellings and Downs curve $\mu_{\rm u}(\gamma)$. Moreover, if the number of sources is large, then these terms make contributions to the pulsar average which are comparable to the first (diagonal) term.
We demonstrate this for the simple “confusion-noise model” constructed in [5]. In this model, each of the $N$ sources of GWs is circularly polarized, and all of them radiate at the same GW frequency $f$, which is picked to be commensurate with the PTA observation interval. From (1cdi), the waveforms are
$h_{n}(t)=A_{n}\exp[i(2\pi ft+\phi_{n})]\,,$ (1cdmaubfbhbo)
where the amplitudes $A_{n}>0$ are real, the GW frequency $f$ is the same fixed multiple of $1/T$ for all $N$ sources, and the phase $\phi_{n}$ of the $n$’th source is a real number in the interval $[0,2\pi)$. For this model, the time averages which appear in (4) are $\overline{h^{*}_{n}h_{m}}=A_{n}A_{m}\exp[i(\phi_{m}-\phi_{n})]$ and $\overline{h^{*}_{n}h^{*}_{m}}=0$. Thus, the pulsar-averaged correlation (4) simplifies to
$\hskip 36.135pt\langle\rho_{ab}\rangle_{ab\in\gamma}=\frac{1}{2}\Bigl{(}\sum_{n}A^{2}_{n}\Bigr{)}\mu_{\rm u}(\gamma)+\frac{1}{2}\sum_{n\neq m}\!A_{n}A_{m}\cos\Delta_{nm}\,\mu(\gamma,\beta_{nm})\,,$
where $\Delta_{nm}\equiv\phi_{n}-\phi_{m}-2\chi^{\phantom{{}^{\prime}}}_{nm}$ is a real phase. Because the “cross-sections” of $\mu(\gamma,\beta)$ for $\beta\neq 0$ do not have exactly the same shape as the HD curve, this means that in any given realization of the universe, the pulsar-averaged correlation will not match the HD curve exactly. (If we could take the correlation curves from many different random realizations of the universe and average them together, then the terms from the second sum would average to zero, and we would obtain exactly the HD curve.)